# Sums related to Euler’s totient function

### v4

@@ 1. Introduction
$\sum_{n=1}^{N}\bigg{(}\frac{a_{n}}{\varphi(a_{n})}\bigg{)}^{s}\leq\textup{exp}(s\ln\ln(s+2)+Cs)N$ (1.1)
@@
From Theorems 1.1 and 1.2 we obtain the following results.
Corollary 1.1.
@@
$\sum_{n\leq x}\Big{(}\frac{f(n)}{\varphi(f(n))}\Big{)}^{s}\leq\textup{exp}(s\ln\ln(s+2)+Cs)x$ (1.3)
@@
Corollary 1.1 extends Corollary 1.1 in [3] which showed the inequality (1.3) but with an upper bound $\textup{exp}(s\ln s+Cs)x$ and without the inequality (1.4).
@@
Corollary 1.2.
Let $s>0$, $a>0$, $b_{1}<\ldots<b_{k}$ be integers. Let $k\geq 2,$ $x\geq 3$, $(\ln x)^{-9/10}\leq\eta\leq 1$, and $|b_{i}|\leq\ln x$ for all $1\leq i\leq k$. We set
@@
1) If $k\geq\ln\ln x$ or $s\leq k$, then
$\sum_{\begin{subarray}{c}|b|\leq\eta\ln x\\ b\,\neq\,b_{1},\ldots,b_{k}\end{subarray}}\Big{(}\frac{a^{k+1}|f(b)|}{\varphi(a^{k+1}|f(b)|)}\Big{)}^{s}\leq\Big{(}c\,\frac{a}{\varphi(a)}\ln k\Big{)}^{s}\eta\ln x.$ (1.5)
2) If $2\leq k<\ln\ln x$ and $s>k$, then
$\sum_{\begin{subarray}{c}|b|\leq\eta\ln x\\ b\,\neq\,b_{1},\ldots,b_{k}\end{subarray}}\Big{(}\frac{a^{k+1}|f(b)|}{\varphi(a^{k+1}|f(b)|)}\Big{)}^{s}\leq\Big{(}c\,\frac{a}{\varphi(a)}\ln s\Big{)}^{s}\eta\ln x.$ (1.6)
@@
Corollary 1.2 extends a result of Maynard ([2, Lemma 8.1]) which showed the same result but with $s=1$ and extends Theorem 1.4 in [3] which showed the inequality (1.6) but with an upper bound
$\Big{(}c\,\frac{a}{\varphi(a)}\ln k\Big{)}^{s}s!\,\eta\ln x.$
@@
Corollary 1.3.
There are absolute positive constants $C$, $c_{1}$, and $c_{2}$ such that for any $x\geq 2$, $s\in\mathbb{N}$, and $t>0$ we have
$\sum_{p\leq x}\Big{(}\frac{p-1}{\varphi(p-1)}\Big{)}^{s}\leq\textup{exp}(s\ln\ln(s+2)+Cs)\pi(x)$ (1.7)
@@
#$\#\big{\{}p\leq x:\frac{p-1}{\varphi(p-1)}>t\big{\}}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))\pi(x).$ (1.8)
@@ 2. Notation
We reserve the letter $p$ for primes. In particular, the sum $\sum_{p\leq K}$ should be interpreted as being over all prime numbers not exceeding $K$. By $\pi(x)$ we denote the number of primes not exceeding $x$. Let #$\#A$ denote the number of elements of a finite set $A$. By $\mathbb{N}$ we denote the set of all positive integers. Let $(a_{1},\ldots,a_{n})$ be the greatest common divisor of integers $a_{1},\ldots,a_{n}$, and $[a_{1},\ldots,a_{n}]$ the least common multiple of integers $a_{1},\ldots,a_{n}$. For real numbers $x,$ $y$ we also use $(x,y)$ to denote the open interval, and $[x,y]$ to denote the closed interval. The usage of the notation will be clear from the context.
@@
We write $\nu(n)$ for the number of distinct primes dividing $n$. By $\mu(n)$ we denote the Möbius mu function, which is defined to be $\mu(n)=(-1)^{\nu(n)}$ if $n$ is square-free, $\mu(n)=0$ otherwise. Let $P^{+}(n)$ denote the greatest prime factor of $n$ (by convention $P^{+}(1)=1$).
@@
The symbol $b|a$ means that $b$ divides $a$. For fixed $a$ the sum $\sum_{b|a}$ and the product $\prod_{b|a}$ should be interpreted as being over all positive divisors of $a$. If $x$ is a real number, then $[x]$ denotes its integral part and $\{x\}=x-[x]$.
@@
$\sum_{n=1}^{N}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}\leq\textup{exp}(s\ln\ln(s+2)+Cs)N,$
@@
$\textup{exp}(s\ln\ln(s+2)+Cs)N$ $\geq\sum_{n=1}^{N}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}\geq\sum_{n\leq N:\,a_{n}/\varphi(a_{n})>t}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}$
@@
#$\#\Big{\{}n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big{\}}\leq\textup{exp}(s\ln\ln(s+2)+Cs-s\ln t)N.$ (3.5)
@@
$\ln(s+2)$ $=te^{-(C+1)}+\ln\Big{(}1+\frac{3-\theta}{\textup{exp}(te^{-(C+1)})}\Big{)}$
@@
$\ln\ln(s+2)$ $=\ln t-(C+1)+\ln\Big{(}1+\frac{e^{(C+1)}R_{1}}{t}\Big{)}$
$=\ln t-(C+1)+R_{2},\qquad 0<R_{2}\leq\frac{3e^{C+1}}{t\,\textup{exp}(te^{-(C+1)})}.$
@@
$s\ln\ln(s+2)+Cs-s\ln t\leq-\textup{exp}(te^{-(C+1)})+\frac{6e^{C+1}}{t}\leq-\textup{exp}(te^{-(C+1)}/2),$ (3.6)
@@
4. Proof of Corollaries 1.1 - 1.3
Proof of Corollary 1.1..
We set $a_{n}=f(n)$. It is clear that $N=[x]$. There exists a positive constant $\tau$ depending only on $f$ such that $f(n)\leq\tau n^{d}$ for any positive integer $n$. We take $M=\tau x^{d}$ and $\alpha=1/2$. We can assume that $x\geq x_{0}$, where $x_{0}>0$ is a large constant depending only on $f$. We have
$y=(\ln M)^{\alpha}=(d\ln x+\ln\tau)^{1/2}\leq(\ln x)^{3/4},$
@@
$n\leq\prod_{p\leq y}p\leq\textup{exp}(2y)\leq\textup{exp}(2(\ln x)^{3/4})\leq\sqrt{x}<[x]=N.$
@@
where $b_{0},\ldots,b_{d}$ are integers with $(b_{0},\ldots,b_{d},m)=1$. Let $\rho(f,m)$ denote the number of solutions of the congruence $f(x)\equiv 0$ (mod $m$). Then
@@
By Theorem 1.2 there exist positive constants $C$, $c_{1}$, and $c_{2}$ (depending only on $d$) such that (1.3) and (1.4) hold. This completes the proof of Corollary 1.1.
Proof of Corollary 1.2..
We can assume that $x\geq x_{0}$, where $x_{0}$ is a large absolute constant. We set
$\Omega=\{b\in\mathbb{Z}:|b|\leq\eta\ln x,b\neq b_{1},\ldots,b_{k}\}.$
@@
Let $b\in\Omega$. Hence, $|b-b_{i}|\leq 2\ln x$ for all $1\leq i\leq k$, and $|f(b)|\leq(2\ln x)^{k}$. Suppose that $k\geq\ln\ln x$. Since $n/\varphi(n)\leq c\ln\ln(n+2)$ for any positive integer $n$, where $c$ is an absolute positive constant, we obtain
$\frac{|f(b)|}{\varphi(|f(b)|)}\leq c(\ln k+\ln\ln(4\ln x))\leq c_{1}\ln k.$
@@
$S\leq\Big{(}\frac{a}{\varphi(a)}\,c_{1}\ln k\Big{)}^{s}\#\Omega\leq\Big{(}c_{2}\,\frac{a}{\varphi(a)}\ln k\Big{)}^{s}\eta\ln x,$
@@
Suppose that $2\leq k<\ln\ln x$. We take $\alpha=1/4$, $M=(2\ln x)^{k}$. Then
$y=(\ln M)^{\alpha}\leq 2(\ln\ln x)^{1/2}.$
@@
$n\leq\prod_{p\leq y}p\leq\textup{exp}(2y)\leq\textup{exp}(4(\ln\ln x)^{1/2})\leq\eta\ln x.$
@@
$\omega(n)=\#\{b\in\Omega:f(b)\equiv 0\text{ (mod $n$)}\}$ $\leq\Big{(}\frac{2\eta\ln x}{n}+1\Big{)}\prod_{p|n}\min(p,k)$
$\leq\frac{3\eta\ln x}{n}\prod_{p|n}\min(p,k).$
@@
$\sum_{b\in\Omega}\Big{(}\frac{|f(b)|}{\varphi(|f(b)|)}\Big{)}^{s}\leq c^{s}(\eta\ln x)\prod_{p\leq y}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}$
$\leq c^{s}(\eta\ln x)\prod_{p}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}$ (4.2)
@@
$S\leq\Big{(}c\,\frac{a}{\varphi(a)}\Big{)}^{s}\eta\ln x\prod_{p}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}.$
Suppose that $s\leq k$ (we recall that $2\leq k<\ln\ln x$). If $p>k$, then $p>s$ and (3.2) holds. Since $g(p)=p/k$, we obtain
@@
$\prod_{p\leq k}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}\leq\textup{exp}\Big{(}\sum_{p\leq k}\frac{s}{p}\Big{)}\leq(c_{2}\ln k)^{s},$
@@
Suppose that $s>k$ and $2\leq k<\ln\ln x$. Since $g(p)\geq p/k>p/s$ for any prime $p$, we obtain
@@
$\prod_{p\leq s}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}\leq\textup{exp}\Big{(}\sum_{p\leq s}\frac{s}{p}\Big{)}\leq(c_{4}\ln s)^{s}.$
The inequality (1.6) is proved. This completes the proof of Corollary 1.2.
Proof of Corollary 1.3..
We set $a_{n}=p_{n}-1$, where $p_{n}$ is the $n$th prime. It is clear that $N=\pi(x)$. We take $\alpha=1/2$, $M=x$. We can assume that $x\geq x_{0}$, where $x_{0}$ is a large absolute constant. Then $y=(\ln x)^{1/2}$. If $n\in\mathcal{D}$, then
$n\leq\prod_{p\leq y}p=\textup{exp}((1+o(1))y)\leq\textup{exp}(2(\ln x)^{1/2})\leq\sqrt{x}.$
By the Brun - Titchmarsh inequality (see, for example, [4, Exercise 9.3.13]) we have
$\omega(n)=\#\{p\leq x:p\equiv 1\text{ {(}mod $n${)}}\}\leq\frac{3x}{\varphi(n)\ln(2x/n)}\leq\gamma\,\frac{\pi(x)}{\varphi(n)},$
where $\gamma$ is an absolute positive constant. Since $\varphi(n)$ is a multiplicative function, $\varphi(p)=p-1\geq 1$ for any prime $p$ (hence, we can take $c_{0}=1$), and
$L=\sum_{p}\frac{1}{p\varphi(p)}=\sum_{p}\frac{1}{p(p-1)}<\infty,$
we can apply Theorem 1.2. Since $\gamma$, $L$, $c_{0}$, and $\alpha$ are absolute positive constants, the constants $C$, $c_{1}$, and $c_{2}$ in (1.7) and (1.8) are absolute and positive. Corollary 1.3 is proved. ∎

### v5

@@ 1. Introduction
$\sum_{n=1}^{N}\bigg{(}\frac{a_{n}}{\varphi(a_{n})}\bigg{)}^{s}\leq\textup{exp}(s\log\log(s+2)+Cs)N$ (1.1)
@@
From Theorems 1.1 and 1.2 we obtain the following result.
Theorem 1.3.
@@
$\sum_{n\leq x}\Big{(}\frac{f(n)}{\varphi(f(n))}\Big{)}^{s}\leq\textup{exp}(s\log\log(s+2)+Cs)x$ (1.3)
@@
Theorem 1.3 extends Corollary 1.1 in [3] which showed the inequality (1.3) but with an upper bound $\textup{exp}(s\log s+Cs)x$ and without the inequality (1.4).
@@
Theorem 1.4.
Let $b_{1},\ldots,b_{k}$ be distinct integers, $a\in\mathbb{N}$, $s\in\mathbb{N}$. Let $k\geq 2,$ $x\geq 3$, $(\ln x)^{-9/10}\leq\eta\leq 1$, and $|b_{i}|\leq\log x$ for all $1\leq i\leq k$. We set
@@
1) If $k\geq\log\log x$ or $s\leq k$, then
$\sum_{\begin{subarray}{c}|b|\leq\eta\log x\\ b\,\neq\,b_{1},\ldots,b_{k}\end{subarray}}\Big{(}\frac{a^{k+1}|f(b)|}{\varphi(a^{k+1}|f(b)|)}\Big{)}^{s}\leq\Big{(}c\,\frac{a}{\varphi(a)}\log k\Big{)}^{s}\eta\log x.$ (1.5)
2) If $2\leq k<\log\log x$ and $s>k$, then
$\sum_{\begin{subarray}{c}|b|\leq\eta\log x\\ b\,\neq\,b_{1},\ldots,b_{k}\end{subarray}}\Big{(}\frac{a^{k+1}|f(b)|}{\varphi(a^{k+1}|f(b)|)}\Big{)}^{s}\leq\Big{(}c\,\frac{a}{\varphi(a)}\log s\Big{)}^{s}\eta\log x.$ (1.6)
@@
Theorem 1.4 extends a result of Maynard ([2, Lemma 8.1]) which showed the same result but with $s=1$ and extends Theorem 1.4 in [3] which showed the inequality (1.6) but with an upper bound
$\Big{(}c\,\frac{a}{\varphi(a)}\log k\Big{)}^{s}s!\,\eta\log x.$
@@
Theorem 1.5.
Let $f(n)=b_{d}n^{d}+\ldots+b_{0}$ be a polynomial with integer coefficients such that $b_{d}>0$, $(b_{d},\ldots,b_{0})=1$, and $f:\mathbb{P}\to\mathbb{N}$. Then there exist positive constants $C$, $c_{1}$, and $c_{2}$ depending only on $f$ such that for any $x\geq 2$, $s\in\mathbb{N}$, and $t>0$ we have
$\sum_{p\leq x}\Big{(}\frac{f(p)}{\varphi(f(p))}\Big{)}^{s}\leq\textup{exp}(s\log\log(s+2)+Cs)\pi(x)$ (1.7)
@@
#$\#\big{\{}p\leq x:\frac{f(p)}{\varphi(f(p))}>t\big{\}}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))\pi(x).$ (1.8)
In particular, we can take $f(x)=x-1$ or $f(x)=x^{2}+1$ in Theorem 1.5.
@@ 2. Notation
We reserve the letter $p$ for primes. In particular, the sum $\sum_{p\leq K}$ should be interpreted as being over all prime numbers not exceeding $K$. We denote the number of primes not exceeding $x$ by $\pi(x)$ and the number of primes not exceeding $x$ that are congruent to $a$ modulo $k$ by $\pi(x;k,a)$. Let #$\#A$ denote the number of elements of a finite set $A$. We write $\mathbb{Z}$ for the set of all integers, $\mathbb{N}$ for the set of all positive integers, $\mathbb{R}$ for the set of all real numbers, and $\mathbb{P}$ for the set of all primes. Let $(a_{1},\ldots,a_{n})$ be the greatest common divisor of integers $a_{1},\ldots,a_{n}$ and $[a_{1},\ldots,a_{n}]$ their least common multiple. For real numbers $x,$ $y$ we also use $(x,y)$ to denote the open interval and $[x,y]$ to denote the closed interval. The usage of the notation will be clear from the context.
@@
We write $\nu(n)$ for the number of distinct primes dividing $n$. By $\mu(n)$ we denote the Möbius function, which is defined to be $\mu(n)=(-1)^{\nu(n)}$ if $n$ is square-free, $\mu(n)=0$ otherwise. Let $P^{+}(n)$ denote the greatest prime factor of $n$ (by convention $P^{+}(1)=1$).
@@
The symbol $b|a$ means that $b$ divides $a$. For fixed $a$ the sum $\sum_{b|a}$ and the product $\prod_{b|a}$ should be interpreted as being over all positive divisors of $a$. By $[x]$ we denote the largest integer not exceeding $x$ and $\{x\}=x-[x]$.
For a polynomial $f(x)$ with integer coefficients, by $\rho(f,m)$ we denote the number of solutions of the congruence $f(x)\equiv 0$ (mod $m$).
@@
$\sum_{n=1}^{N}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}\leq\textup{exp}(s\log\log(s+2)+Cs)N,$
@@
$\textup{exp}(s\log\log(s+2)+Cs)N$ $\geq\sum_{n=1}^{N}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}\geq\sum_{n\leq N:\,a_{n}/\varphi(a_{n})>t}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}$
@@
#$\#\Big{\{}n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big{\}}\leq\textup{exp}(s\log\log(s+2)+Cs-s\log t)N.$ (3.5)
@@
$\log(s+2)$ $=te^{-(C+1)}+\log\Big{(}1+\frac{3-\theta}{\textup{exp}(te^{-(C+1)})}\Big{)}$
@@
$\log\log(s+2)$ $=\log t-(C+1)+\log\Big{(}1+\frac{e^{(C+1)}R_{1}}{t}\Big{)}$
$=\log t-(C+1)+R_{2},\qquad 0<R_{2}\leq\frac{3e^{C+1}}{t\,\textup{exp}(te^{-(C+1)})}.$
@@
$s\log\log(s+2)+Cs-s\log t\leq-\textup{exp}(te^{-(C+1)})+\frac{6e^{C+1}}{t}\leq-\textup{exp}(te^{-(C+1)}/2),$ (3.6)
@@
4. Proof of Theorems 1.3 and 1.4
Proof of Theorem 1.3..
We set $a_{n}=f(n)$. It is clear that $N=[x]$. There exists a positive constant $\tau$ depending only on $f$ such that $f(n)\leq\tau n^{d}$ for any positive integer $n$. We take $M=\tau x^{d}$ and $\alpha=1/2$. We assume that $x\geq x_{0}$, where $x_{0}>0$ is a large constant depending only on $f$. We have
$y=(\log M)^{\alpha}=(d\log x+\log\tau)^{1/2}\leq(\log x)^{3/4},$ (4.1)
@@
$n\leq\prod_{p\leq y}p\leq\textup{exp}(2y)\leq\textup{exp}(2(\log x)^{3/4})\leq\sqrt{x}<[x]=N.$ (4.2)
@@
where $b_{0},\ldots,b_{d}$ are integers with $(b_{0},\ldots,b_{d},m)=1$. Then
@@
By Theorem 1.2 there exist positive constants $C$, $c_{1}$, and $c_{2}$ (depending only on $d$) such that (1.3) and (1.4) hold.
Suppose that $1\leq x<x_{0}$. For any $n\leq x$ we have
$\frac{f(n)}{\varphi(f(n))}\leq f(n)\leq\tau n^{d}\leq\tau x_{0}(f)^{d}=A,$
where $A=A(f)>0$ is a constant depending only on $f$. Hence,
$\sum_{n\leq x}\Big{(}\frac{f(n)}{\varphi(f(n))}\Big{)}^{s}\leq A^{s}x\leq\textup{exp}(s\log\log(s+2)+Bs)x,$
where $B=\log A$ is a positive constant depending only on $f$.
We set
$\lambda(t)=\#\Big{\{}n\leq x:\frac{f(n)}{\varphi(f(n))}>t\Big{\}}.$
If $0<t\leq A$, then
$\lambda(t)\leq[x]\leq x_{0}\leq x_{0}\frac{\textup{exp}(\textup{exp}(c_{2}A))}{\textup{exp}(\textup{exp}(c_{2}t))}=\frac{b_{1}}{\textup{exp}(\textup{exp}(c_{2}t))}\leq\frac{b_{1}}{\textup{exp}(\textup{exp}(c_{2}t))}x,$
where $b_{1}=x_{0}\,\textup{exp}(\textup{exp}(c_{2}A))$ is a positive constant depending only on $f$. If $t>A$, then
$\lambda(t)=0\leq\frac{b_{1}}{\textup{exp}(\textup{exp}(c_{2}t))}x.$
We put
$\widetilde{C}=\max(C,B),\qquad\widetilde{c}_{1}=\max(c_{1},b_{1}),\qquad\widetilde{c}_{2}=c_{2}.$
Then $\widetilde{C}$, $\widetilde{c_{1}}$ and $\widetilde{c}_{2}$ are positive constants depending only on $f$, and for any $x\geq 1$, $s\in\mathbb{N}$, and $t>0$ we have
$\sum_{n\leq x}\Big{(}\frac{f(n)}{\varphi(f(n))}\Big{)}^{s}\leq\textup{exp}(s\log\log(s+2)+\widetilde{C}s)x$
and
#$\#\Big{\{}n\leq x:\frac{f(n)}{\varphi(f(n))}>t\Big{\}}\leq\frac{\widetilde{c}_{1}}{\textup{exp}(\textup{exp}(\widetilde{c}_{2}t))}x.$
Theorem 1.3 is proved.
Proof of Theorem 1.4..
We assume that $x\geq x_{0}$, where $x_{0}$ is a large absolute constant. We set
$\Omega=\{b\in\mathbb{Z}:|b|\leq\eta\log x,b\neq b_{1},\ldots,b_{k}\}.$
@@
Let $b\in\Omega$. Hence, $|b-b_{i}|\leq 2\log x$ for all $1\leq i\leq k$, and $|f(b)|\leq(2\log x)^{k}$. Suppose that $k\geq\log\log x$. Since $n/\varphi(n)\leq c\log\log(n+2)$ for any positive integer $n$, where $c$ is an absolute positive constant, we obtain
$\frac{|f(b)|}{\varphi(|f(b)|)}\leq c(\log k+\log\log(4\log x))\leq c_{1}\log k.$
@@
$S\leq\Big{(}\frac{a}{\varphi(a)}\,c_{1}\log k\Big{)}^{s}\#\Omega\leq\Big{(}c_{2}\,\frac{a}{\varphi(a)}\log k\Big{)}^{s}\eta\log x,$
@@
Suppose that $2\leq k<\log\log x$. We take $\alpha=1/4$, $M=(2\log x)^{k}$. Then
$y=(\ln M)^{\alpha}\leq 2(\log\log x)^{1/2}.$
@@
$n\leq\prod_{p\leq y}p\leq\textup{exp}(2y)\leq\textup{exp}(4(\log\log x)^{1/2})\leq\eta\log x.$
@@
$\omega(n)=\#\{b\in\Omega:f(b)\equiv 0\text{ (mod $n$)}\}$ $\leq\Big{(}\frac{2\eta\log x}{n}+1\Big{)}\prod_{p|n}\min(p,k)$
$\leq\frac{3\eta\log x}{n}\prod_{p|n}\min(p,k).$
@@
$\sum_{b\in\Omega}\Big{(}\frac{|f(b)|}{\varphi(|f(b)|)}\Big{)}^{s}\leq c^{s}(\eta\log x)\prod_{p\leq y}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}$
$\leq c^{s}(\eta\log x)\prod_{p}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}$ (4.4)
@@
$S\leq\Big{(}c\,\frac{a}{\varphi(a)}\Big{)}^{s}\eta\log x\prod_{p}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}.$
Suppose that $s\leq k$ (we recall that $2\leq k<\log\log x$). If $p>k$, then $p>s$ and (3.2) holds. Since $g(p)=p/k$, we obtain
@@
$\prod_{p\leq k}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}\leq\textup{exp}\Big{(}\sum_{p\leq k}\frac{s}{p}\Big{)}\leq(c_{2}\log k)^{s},$
@@
Suppose that $s>k$ and $2\leq k<\log\log x$. Since $g(p)\geq p/k>p/s$ for any prime $p$, we obtain
@@
$\prod_{p\leq s}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}\leq\textup{exp}\Big{(}\sum_{p\leq s}\frac{s}{p}\Big{)}\leq(c_{4}\log s)^{s}.$
The inequality (1.6) is proved.
Suppose that $3\leq x<x_{0}$. For any $b\in\Omega$ we have $|f(b)|\leq(2\log x)^{k}\leq A_{0}^{k}$, where $A_{0}=2\log x_{0}$, and
$\frac{|f(b)|}{\varphi(|f(b)|)}\leq c\log\log(|f(b)|+2)\leq A_{1}\log k,$
where $A_{1}>0$ is an absolute constant. Since
#$\#\Omega\leq 2\eta\log x+1\leq 2\log x_{0}+1$
and
$\eta\log x\geq\frac{1}{(\log x)^{1/10}}\geq\frac{1}{(\log x_{0})^{1/10}},$
by (4.3) we have
$S\leq\Big{(}A_{1}\frac{a}{\varphi(a)}\log k\Big{)}^{s}\#\Omega\leq\Big{(}A_{2}\frac{a}{\varphi(a)}\log k\Big{)}^{s}\eta\log x,$
where $A_{2}>0$ is an absolute constant. This completes the proof of Theorem 1.4.
5. Proof of Theorem 1.5
Proof.
We set $a_{p}=f(p)$. It is clear that $N=\pi(x)$. There exists a positive constant $\tau$ depending only on $f$ such that $f(n)\leq\tau n^{d}$ for any positive integer $n$. We take $M=\tau x^{d}$ and $\alpha=1/2$. We can assume that $x\geq x_{0}$, where $x_{0}>0$ is a large constant depending only on $f$. By (4.1) and (4.2) we have $y=(\log M)^{\alpha}\leq(\log x)^{3/4}$ and $n\leq\sqrt{x}$ for any $n\in\mathcal{D}$.
We need the following result (see, for example, [4, Exercise 9.3.13]).
Lemma 5.1 (The Brun - Titchmarsh inequality).
For any $\varepsilon>0$ there exists a constant $x_{0}(\varepsilon)>0$ depending only on $\varepsilon$ such that for any $a\in\mathbb{Z}$, $k\in\mathbb{N}$, and $x\in\mathbb{R}$ with $(a,k)=1$, $k\leq x$ and $x\geq x_{0}(\varepsilon)$ we have
$\pi(x;k,a)\leq\frac{(2+\varepsilon)x}{\varphi(k)\log(2x/k)}.$
Fix $n\in\mathcal{D}$, and let $\alpha_{1},\ldots,\alpha_{l}$ be all solutions of the congruence $f(x)\equiv 0$ (mod $n$). For any prime $p|n$, by Lagrange’s theorem we have $\rho(f,p)\leq\min(p,d)$. Hence,
$l=\rho(f,n)=\prod_{p|n}\rho(f,p)\leq\prod_{p|n}\min(p,d).$ (5.1)
We have
$\omega(n)=\#\{p\leq x:f(p)\equiv 0\text{ (mod $n$)}=\sum_{i=1}^{l}\pi(x;n,\alpha_{i}).$ (5.2)
Let $1\leq i\leq s$, and suppose that $(\alpha_{i},n)=1$. We have
$\log\Big{(}\frac{2x}{n}\Big{)}\geq\log(2\sqrt{x})\geq\frac{1}{2}\log x.$
Applying Lemma 5.1 with $\varepsilon=1/2$, we obtain
$\pi(x;n,\alpha_{i})\leq\frac{5x}{\varphi(n)\log x},$ (5.3)
if $x_{0}$ is large enough.
Suppose that $(\alpha_{i},n)>1$. Hence,
$\pi(x;n,\alpha_{i})\leq 1\leq\frac{5x}{\varphi(n)\log x},$
since
$\frac{5x}{\varphi(n)\log x}\geq\frac{5x}{n\log x}\geq\frac{5\sqrt{x}}{\log x}>1,$
if $x_{0}$ is large enough. We obtain that the inequality (5.3) holds in both cases $(\alpha_{i},n)=1$ and $(\alpha_{i},n)>1$. From (5.1) – (5.3) we get
$\omega(n)\leq\frac{5xl}{\varphi(n)\log x}\leq\frac{\gamma\pi(x)}{\varphi(n)}\prod_{p|n}\min(p,d),$
where $\gamma>0$ is an absolute constant.
We set $g(1)=1,$
$g(p^{\beta})=\begin{cases}(p-1)/\min(p,d),&\text{if $\beta=1$;}\\ 1,&\text{if $\beta\geq 2$};\end{cases}$
$\omega(n)\leq\frac{\gamma\pi(x)}{g(n)}$
for any $n\in\mathcal{D}$.
Since
$g(p)=\frac{p-1}{\min(p,d)}\geq\frac{p-1}{p}\geq\frac{1}{2}$
for any prime $p$ and
$L=\sum_{p}\frac{1}{pg(p)}=\sum_{p\leq d}\frac{1}{p-1}+\sum_{p>d}\frac{d}{p(p-1)}<\infty,$
by Theorem 1.2 there exist positive constants $C$, $c_{1}$ and $c_{2}$ depending only on $d$ such that the inequalities (1.7) and (1.8) hold.
By arguing as in the end of the proof of Theorem 1.3, we find that there exist positive constants $\widetilde{C}$, $\widetilde{c}_{1}$, and $\widetilde{c}_{2}$ depending only on $f$ such that for any $x\geq 2$, $s\in\mathbb{N}$, and $t>0$ we have
$\sum_{p\leq x}\Big{(}\frac{f(p)}{\varphi(f(p))}\Big{)}^{s}\leq\textup{exp}(s\log\log(s+2)+\widetilde{C}s)\pi(x)$
and
#$\#\Big{\{}p\leq x:\frac{f(p)}{\varphi(f(p))}>t\Big{\}}\leq\frac{\widetilde{c}_{1}}{\textup{exp}(\textup{exp}(\widetilde{c}_{2}t))}\pi(x).$
Theorem 1.5 is proved.